# Bayesian Constraints on Pre-Equilibrium Jet Quenching and Predictions for Oxygen Collisions

### v1

@@ II Framework
$\begin{split}\sigma^{AA}(p_{T},R)&=\sum_{i=q,g}\int d\varepsilon P_{i}(\varepsilon,R,p_{T})\tilde{\sigma}^{pp}_{i}(p_{T}+\varepsilon,R)\,.\end{split}$ (1)
@@
Energy is lost predominantly due to medium-induced emissions escaping the jet cone $R$, as encoded in $P_{i}(\varepsilon,R,p_{T})$. The emission rate is computed using the Improved Opacity Expansion in a static medium, which captures both single hard and multiple soft scatterings, and also includes finite path length corrections [52, 53]. Medium properties affect these rates through LO Hard Thermal Loop results, namely the (bare) jet quenching parameter $\hat{q}_{0}=g^{2}_{\rm med}N_{c}m_{D}^{2}T/(4\pi)$, and the Debye screening mass $m_{D}^{2}=3g_{\rm med}^{2}T^{2}/2$, where $T$ is the local temperature of the hydrodynamic medium. Here, $g_{\rm med}$ is the (fixed, i.e. non-running) strong coupling for jet-medium interactions, representing the first of the two parameters we will estimate using Bayesian inference. We assume factorization between induced collinear emissions and transverse momentum broadening: semi-hard emissions undergo Gaussian broadening, while hard emissions inherit their transverse momentum from the Coulomb tail of the scattering potential. Since the softest emissions thermalize rapidly [54], they should be considered part of the medium and contribute to hydrodynamic wakes [55, 56], a manifestation of medium response to the jet passage. We approximate these physics by assuming that the thermalized energy is distributed uniformly around the jet hemisphere. Elastic energy loss is also included and constrained via the Einstein relation $\hat{e}=\hat{q}/(4T)$, and is assumed to thermalize in a similar fashion. Here, $\hat{q}$ accounts for the logarithmic corrections of $\hat{q}_{0}$ (see later).
@@
Energetic jets consist of multiple partons produced by the high-virtuality DGLAP evolution that occurs at the early stages of the in-medium jet propagation [57]. These partons represent an additional source of energy loss. We account for this by resumming these multiple sources via a factorization between early vacuum-like and medium-induced emissions [58]. Importantly, we also include color-coherence, and -resolution [59, 60, 61] effects (referred to as coherent energy loss) by incorporating the non-local angular scale below which the medium cannot resolve individual partons, $\theta_{c}\sim 1/\sqrt{\hat{q}L^{3}}$, with $L$ being the traversed medium length. Detailed formulas are provided in [44].
@@
Jet propagation occurs within an event-by-event fluctuating 2+1D hydrodynamic background. We employ hydro profiles from the comprehensive multi-stage framework IP-Glasma+JIMWLK+MUSIC+UrQMD, which successfully describes a wide range of bulk observables in heavy-ion collisions [62]. This state-of-the-art framework evolves the bulk starting from the initial energy density deposited at specific hotspots determined by the IPSat [63]. Jet probes are placed precisely onto these hotspots, whose number is proportional to the number of binary collisions $N_{\rm coll}$, and their initial orientation is uniformly distributed in the azimuthal angle $\phi$. Straight jet trajectories are sampled from their production point until they exit the deconfined medium at $T<T_{c}$, with $T_{c}$ marks the crossover transition to the hadron gas phase. The medium and jet-medium interaction quantities needed for the computation of radiative and elastic energy loss are then obtained by averaging along each jet trajectory, as extracted from the hydrodynamic profiles. We neglect further quenching below $T_{c}$. For more details on the embedding of our jet quenching framework in the realistic heavy-ion environment, see Ref. [44], where event-averaged hydro profiles were used instead. Event classification into centrality classes is based on the final charged particle multiplicity of the bulk, from which one also extracts the soft event-plane angles $\Psi_{n}$, necessary for computing the jet flow coefficients. Thus, our framework predicts the quenched jet spectrum for various parameters $\sigma^{AA\%}(p_{T},\eta,\phi,R)$.
@@
The hydrodynamic description of the bulk begins at a finite time $\tau_{\rm hyd}=0.4$ fm. The initial condition of this nearly ideal fluid is reached through the Glasma evolution of saturated gluon fields (for reviews on this approach to thermalization, see [64, 65]). Quenching effects have traditionally been neglected before this time, largely due to the absence of reliable energy-loss calculations in the non-equilibrated phase. While an earlier start could be mimicked by adjusting the jet-medium coupling, combined observables such as high-$p_{T}$ $R_{AA}$ and $v_{2}$ are known to be sensitive to the actual value of the quenching onset time [66, 67]. The need to address jet quenching in the earliest stages has been emphasized in recent works where the pre-hydro phase is modeled either by Glasma [68, 69] or by QCD effective kinetic theory (EKT) [70, 71]. Despite differences in approach, these studies consistently conclude that the broadening per unit length in the short pre-equilibrium phase is considerably larger than in the equilibrated hydro phase.
@@
In the present work, we extend our energy loss framework to the pre-equilibrium stage as follows. The out-of-equilibrium evolution of the local energy density has been extensively studied in EKT [72, 73, 74, 75, 76, 77], and it is well captured by the hydrodynamic attractor. We use the attractor to extrapolate the effective local temperature for times $\tau$ earlier than $\tau_{\rm hyd}$ as [78, 79]:
@@
with $\eta/s=0.12$, where $T(x)$ is the local initial temperature. We assume that our energy-loss model can be applied without any additional modification by using these extrapolated pre-equilibrium temperatures. Before $\tau_{\rm min}$, no quenching is applied, and we take this time as the second parameter of our Bayesian inference.
Finally, fluid velocity fields also evolve during the pre-equilibrium phase, as initial spatial anisotropies begin to generate flow [65]. This non-equilibrium evolution is well approximated by the universal pre-flow extracted from EKT [80]. We incorporate it by scaling the transverse velocity linearly with time:
@@ III Results
$\begin{split}R^{\rm jet}_{AA}=&\sigma^{AA\%}(p_{T},R)/\sigma^{pp}(p_{T},R)\,,\\ v^{\rm jet}_{n}=&\frac{\int d\eta d\phi\cos(n(\phi-\Psi_{n}))\sigma^{AA\%}(p_{T},\eta,\phi,R)}{\sigma^{AA}(p_{T},R)}\,.\end{split}$ (5)
where $\sigma(p_{T},R)=\int d\eta d\phi\sigma(p_{T},\eta,\phi,R)$ with the appropriate experimental cuts. We use the event plane method to compute $v_{n}^{\rm jet}$, following the procedure in the ATLAS measurement to which we compare our model [81]. We evaluate the hadron spectrum as
$\sigma(p^{h}_{T},\eta,\phi)=\sum_{i=q,g}\int^{1}_{x}\frac{{\rm d}z}{z}\sigma_{i}(\tfrac{p^{h}_{T}}{z},R,\eta,\phi)\,D_{i}^{h}(z,\mu_{F}^{2})\,$, (6)
@@
We employ Bayesian inference, implemented via the JETSCAPE/STAT toolkit [101, 102, 103], to constrain the jet-medium strong coupling $g_{\rm med}$ and the initial quenching time in the pre-equilibrium phase $\tau_{\rm min}$. We used flat priors in the intervals $1.2\leq g_{\rm med}\leq 1.7$, and $0.01\leq\tau_{\rm min}\,[{\rm fm}]\leq 0.4$. Our data points include $R^{\rm jet}_{AA}$ at LHC energies for different jet-cones and centrality selections [104, 105, 106], as well as jet $v_{2}^{\rm jet}$ [81], up to 40-50% centrality. The inference incorporates statistical and systematic uncertainties of the data, and we estimate the full correlation matrix by employing a finite correlation length of 0.2 [102, 107]. Luminosity and centrality uncertainties are added to the systematics. We neglect inherent model uncertainties in the inference such as scale and nPDF uncertainties, and we will present them separately for the OO predictions.
@@
Figure. 1 shows the parameter posterior for two different scenarios. Fitting only $R^{\rm jet}_{AA}$ (blue) results in little constraint on $\tau_{\rm min}$. This is expected, as an earlier starting time can be compensated by a relatively smaller medium coupling to produce the same energy loss. Including $v_{2}^{\rm jet}$ constrains $\tau_{\rm min}\approx 0.24$ fm (red). This inference illustrates the role of pre-equilibrium effects in $v^{\rm jet}_{2}$.
@@
Before moving to our predictions for oxygen collisions, Fig. 4 in App. A shows the posterior predictive distributions. We observe excellent overall agreement with all data. Small tensions can be found in $v_{2}^{\rm jet}$ at low $p_{T}$, which motivates further improvements of the energy loss framework. For completeness, we also show inference results on $g_{\rm med}$ with two limiting scenarios: early pre-equilibrium with fixed $\tau_{\min}=0.08$ fm, and no pre-equilibrium with $\tau_{\min}=0.4$ fm. The best fit (whose parameters are shown in Fig. 1) outperforms these two scenarios, although differences are not very significant. The extracted bare jet transport coefficient is approximately $\hat{q}_{0}/T^{3}\approx 1.7$, while after logarithmic corrections the effective value is $\hat{q}/T^{3}=\hat{q}_{0}/T^{3}\log(Q_{s}^{2}/\mu_{*}^{2})\approx 7$, where $Q_{s}^{2}=\hat{q}_{0}L$, and $\mu_{*}^{2}=m_{D}^{2}e^{-2+2\gamma_{E}}/4$. We note that this extracted $\hat{q}$ is consistent with other recent extractions using Bayesian inference in the JETSCAPE framework [103].
@@
Figure 5 in App. A provides predictions for jets at RHIC energies [108] and charged hadrons at LHC energies [109, 110, 111, 112] using the extracted parameters. Even though our framework was designed for jets, it also provides a good description of the charged hadron $R_{AA}$ without including them in the inference. We also see, however, that our framework underestimates hadron elliptic flow at low $p_{T}$. This failure is consistent with previous findings using pQCD energy loss [66, 67]. It has been pointed out that a rather late quenching time of around $\sim 1$ fm would resolve the disagreement [67]. Alternatively, the absence of this issue for jets signals the need for model improvements which could be more relevant for (mid-$p_{T}$) hadrons than for jets. Again, for completeness, we show predictions with the previous two limiting scenarios (early pre-eq. and no pre-eq. energy loss), resulting as before in subtle differences.
@@
Figure 3 shows the inclusive jet and charged hadron elliptic flow. There is relatively smaller sensitivity to the nPDF, as $v_{2}$ is an observable that is differential in the transverse plane, using solely the medium spectra, so the impact of the nPDF largely cancels in Eq. (5). Semi-inclusive observables can present analogous advantages [115]. We observe that, in contrast to large collision systems, the elliptic flow coefficient $v_{2}$ increases with centrality. This is a consequence of the more prominent role played by the fluctuations of the initial nuclear shape in small collision systems [117, 126, 119]. Jet distributions are not shown for very low $p_{T}$, where our framework is not expected to apply.
Due to the relatively small path lengths involved in OO, the typical value of the coherence angle is relatively large, $\theta_{c}\sim 0.5$, and so $R=0.4$ jets are unresolved and lose energy as single color charges. Therefore, hadron and jet suppression in our framework can be understood by the mere $p_{T}$ shift resulting from hadronization. In contrast, in larger systems (with smaller values of $\theta_{c}$) both $R_{AA}$ and $v_{2}$ receive additional contributions due to resolved jet substructure fluctuations [127, 44].
@@ IV Conclusions
In this work, we have integrated our semi-analytical jet quenching model with a state-of-the-art description of heavy-ion collisions, IP-Glasma+JIMWLK+MUSIC+UrQMD. We incorporated energy loss in the pre-equilibrium phase via the hydrodynamic attractor for the first time. Our framework provides a simultaneous description of jet $R_{AA}$ and $v_{2}$ in agreement with data in $AA$ collisions, offering a strong validation of the approach. We show that a joint description of these observables places constraints on the onset time of jet-medium interactions, which is around 0.2 fm, deep in the pre-equilibrium phase. Then, we applied these parameters to predict jet and hadron quenching in other collision systems. We saw that our framework successfully predicts hadron suppression but underestimates hadron elliptic flow towards lower $p_{T}$. Previous studies have reported similar results, suggesting that a delayed onset of quenching of about 1 fm could account for the hadron experimental data. Our success in describing jets, however, would rather suggest the need to account for additional medium effects whose importance could grow at lower $p_{T}$.
@@
Our constrained framework has been used to make realistic predictions for the intermediate-sized OO collisions. We obtained sizable energy loss both for jets and charged hadrons, in quantitative agreement with the preliminary results from CMS [128]. Even though we have shown that pre-equilibrium effects play a moderate role in our model, we anticipate that including forthcoming data on high-$p_{T}$ flow coefficients will result in further constraints. Furthermore, in this work we have provided predictions for elliptic anisotropies and different centrality selections, both for hadrons and jets. Due to the small path lengths in OO, unlike in PbPb or AuAu, jet constituents are barely resolved (jets are fully coherent), and thus we obtain similar energy loss for hadrons and jets (modulo a $p_{T}$ shift in fragmentation). This outcome constitutes a prediction of our framework which can be confronted with future experimental results on jet suppression in light ions.
@@
As an outlook, our jet quenching framework could benefit from an improved description of medium-induced emissions, elastic scatterings, and medium response in several ways. It is important to note that, at very early times, local color fields are saturated and highly anisotropic, a scenario which our energy-loss framework, built on local equilibrium, cannot accommodate. For these initial stages, it would be desirable to incorporate recent developments in energy-loss calculations for anisotropic media [129, 130, 131, 132, 133, 134], correlated color fields [135], and out of-equilibrium plasma [136, 137], so as to extend quenching for those earliest times where the attractor solution is not meant to apply.

### v2

@@ II Framework
$split\sigma^{AA}(p_{T},R)&=\sum_{i=q,g}\int d\varepsilon P_{i}(\varepsilon,R,p_{T})\tilde{\sigma}^{pp}_{i}(p_{T}+\varepsilon,R)\,.$ (1)
@@
Energy is lost predominantly due to medium-induced emissions escaping the jet cone $R$, as encoded in $P_{i}(\varepsilon,R,p_{T})$. The emission rate is computed using the Improved Opacity Expansion in a static medium, which captures both single hard and multiple soft scatterings, and also includes finite path length corrections [52, 53] [^1] The static medium rates can mimic the rates of expanding medium in the multiple soft scattering limit [54, 55].. Medium properties affect these rates through LO Hard Thermal Loop results, namely the (bare) jet quenching parameter $\hat{q}_{0}=g^{2}_{\rm med}N_{c}m_{D}^{2}T/(4\pi)$, and the Debye screening mass $m_{D}^{2}=3g_{\rm med}^{2}T^{2}/2$, where $T$ is the local temperature of the hydrodynamic medium. Here, $g_{\rm med}$ is the (fixed, i.e. non-running) strong coupling for jet-medium interactions, representing the first of the two parameters we will estimate using Bayesian inference. We assume factorization between induced collinear emissions and transverse momentum broadening: semi-hard emissions undergo Gaussian broadening, while hard emissions inherit their transverse momentum from the Coulomb tail of the scattering potential. Since the softest emissions thermalize rapidly [56], they should be considered part of the medium and contribute to hydrodynamic wakes [57, 58], a manifestation of medium response to the jet passage. We approximate these physics by assuming that the thermalized energy is distributed uniformly around the jet hemisphere. Elastic energy loss is also included and constrained via the Einstein relation $\hat{e}=\hat{q}/(4T)$, and is assumed to thermalize in a similar fashion. Here, $\hat{q}$ accounts for the logarithmic corrections of $\hat{q}_{0}$ (see later).
@@
Energetic jets consist of multiple partons produced by the high-virtuality DGLAP evolution that occurs at the early stages of the in-medium jet propagation [59]. These partons represent an additional source of energy loss. We account for this by resumming these multiple sources via a factorization between early vacuum-like and medium-induced emissions [60]. Importantly, we also include color-coherence, and -resolution [61, 62, 63] effects (referred to as coherent energy loss) by incorporating the non-local angular scale below which the medium cannot resolve individual partons, $\theta_{c}\approx 1/\sqrt{\hat{q}L^{3}}$, with $L$ being the traversed medium length. Detailed formulas are provided in [44].
@@
Jet propagation occurs within an event-by-event fluctuating 2+1D hydrodynamic background. We employ hydro profiles from the comprehensive multi-stage framework IP-Glasma+JIMWLK+MUSIC+UrQMD, which successfully describes a wide range of bulk observables in heavy-ion collisions [64]. This state-of-the-art framework evolves the bulk starting from the initial energy density deposited at specific hotspots determined by the IPSat [65]. Jet probes are placed precisely onto these hotspots, whose number is proportional to the number of binary collisions $N_{\rm coll}$, and their initial orientation is uniformly distributed in the azimuthal angle $\phi$. Straight jet trajectories are sampled from their production point until they exit the deconfined medium at $T<T_{c}$, with $T_{c}$ marks the crossover transition to the hadron gas phase. The medium and jet-medium interaction quantities needed for the computation of radiative and elastic energy loss are then obtained by averaging along each jet trajectory, as extracted from the hydrodynamic profiles. For example, we evaluate the Debye mass as $m_{D}^{2}\propto(\int{\rm d}x_{F}T^{2}(x))/L$ for a given trajectory, where $x_{F}$ is the distance traversed in the fluid rest frame, thereby accounting for different flow velocities, and $L$ is the integral of ${\rm d}x_{F}$. The expressions for other relevant quantities can be found in Eq. (29) in [44]. We neglect further quenching below the freeze-out temperature $T_{c}$. For more details on the embedding of our jet quenching framework in the realistic heavy-ion environment, see Ref. [44], where event-averaged hydro profiles were used instead. Event classification into centrality classes is based on the final charged particle multiplicity of the bulk, from which one also extracts the soft event-plane angles $\Psi_{n}$, necessary for computing the jet flow coefficients. Thus, our framework predicts the quenched jet spectrum for various parameters $\sigma^{AA\%}(p_{T},\eta,\phi,R)$.
@@
The hydrodynamic description of the bulk begins at a finite time $\tau_{\rm hyd}=0.4$ fm. This parameter was fixed when the IP-Glasma+JIMWLK+MUSIC+UrQMD framework was tuned to LHC and RHIC data. The initial condition of this nearly ideal fluid is reached through the Glasma evolution of saturated gluon fields. The Glasma evolution should be followed by a kinetic evolution that drives the system toward hydrodynamics; however, this is not the setup used in the current bulk evolution framework. For reviews of this approach to thermalization, see [66, 67]. Quenching effects have traditionally been neglected before $\tau_{\rm hyd}$, largely due to the absence of reliable energy-loss calculations in the non-equilibrated phases. While an earlier start could be mimicked by adjusting the jet-medium coupling, combined observables such as high-$p_{T}$ $R_{AA}$ and $v_{2}$ are known to be sensitive to the actual value of the quenching onset time [68, 69]. The need to address jet quenching in the earliest stages has been emphasized in recent works where the pre-hydro phase is modeled either by Glasma [70, 71] or by QCD effective kinetic theory (EKT) [72, 73]. Despite differences in approach, these studies consistently conclude that the broadening per unit length in the short pre-equilibrium phase is considerably larger than in the equilibrated hydro phase.
@@
In the present work, we extend our energy loss framework to the kinetic pre-equilibrium stage as follows. The out-of-equilibrium evolution of the local energy density has been extensively studied in EKT [74, 75, 76, 77, 78, 79], and it is well captured by the hydrodynamic attractor. We use the attractor to extrapolate the effective local temperature for times $\tau$ earlier than $\tau_{\rm hyd}$ as [80, 81]:
@@
with $\eta/s=0.12$, where $T(x)$ is the local initial temperature. We assume that our energy-loss model can be applied without any additional modification by using these extrapolated pre-equilibrium temperatures. Before $\tau_{\rm min}$ , and for the Glasma phase, no quenching is applied, and we take this time as the second parameter of our Bayesian inference.
Finally, fluid velocity fields also evolve during the kinetic pre-equilibrium phase, as initial spatial anisotropies begin to generate flow [67]. This non-equilibrium evolution is well approximated by the universal pre-flow extracted from EKT [82]. We incorporate it by scaling the transverse velocity linearly with time:
@@ III Results
$split R^{\rm jet}_{AA}=&\sigma^{AA\%}(p_{T},R)/\sigma^{pp}(p_{T},R)\,,\\ v^{\rm jet}_{n}=&\frac{\int d\eta d\phi\cos(n(\phi-\Psi_{n}))\sigma^{AA\%}(p_{T},\eta,\phi,R)}{\sigma^{AA}(p_{T},R)}\,.$ (5)
where $\sigma(p_{T},R)=\int d\eta d\phi\sigma(p_{T},\eta,\phi,R)$ with the appropriate experimental cuts. We use the event plane method to compute $v_{n}^{\rm jet}$, following the procedure in the ATLAS measurement to which we compare our model [83]. [^2] We also compared our results using the event plane method against the more experimentally common scalar product method, in which one correlates the jet momenta with the soft sector (see e.g. [84]), and found very little change in $v_{2}$, below $4\%$, consistent with what was shown by ATLAS [85]. We evaluate the hadron spectrum as
[…]
@@
We employ Bayesian inference, implemented via the JETSCAPE/STAT toolkit [104, 105, 106], to constrain the jet-medium strong coupling $g_{\rm med}$ and the initial quenching time in the pre-equilibrium phase $\tau_{\rm min}$. We used flat priors in the intervals $1.2\leq g_{\rm med}\leq 1.7$, and $0.01\leq\tau_{\rm min}\,[{\rm fm}]\leq 0.4$. We do not consider the scenario when $\tau_{\min}>\tau_{\rm hyd}$ as we believe delaying quenching beyond hydrodynamics is unphysical. Our data points include $R^{\rm jet}_{AA}$ at LHC energies for different jet-cones and centrality selections [107, 108, 109], as well as jet $v_{2}^{\rm jet}$ [83], up to 40-50% centrality. The inference incorporates statistical and systematic uncertainties of the data, and we estimate the full correlation matrix by employing a finite correlation length of 0.2 [105, 110]. Luminosity and centrality uncertainties are added to the systematics. We neglect inherent model uncertainties in the inference such as scale and nPDF uncertainties, and we will present them separately for the OO predictions.
@@
Figure. 1 shows the parameter posterior for two different scenarios. Fitting only $R^{\rm jet}_{AA}$ (blue) results in little constraint on $\tau_{\rm min}$. This is expected, as an earlier starting time can be compensated by a relatively smaller medium coupling to produce the same energy loss. Including $v_{2}^{\rm jet}$ constrains $\tau_{\rm min}\approx 0.24$ fm (red). This inference illustrates the role of pre-equilibrium effects in $v^{\rm jet}_{2}$. [^3] It would be interesting to include $\tau_{\rm hyd}$ in the inference, since it could affect our results. This would require the simultaneous tuning of the soft and hard sectors, a task that we defer to future work.
@@
Before moving to our predictions for oxygen collisions, Figs. 4-5 in App. A show the posterior predictive distributions. We observe excellent overall agreement with all data. Small tensions can be found in $v_{2}^{\rm jet}$ at low $p_{T}$, which motivates further improvements of the energy loss framework. For completeness, Figs. 4-5 in App. A also show two additional inference results on $g_{\rm med}$ with two limiting scenarios: early pre-equilibrium with fixed $\tau_{\min}=0.08$ fm, and no pre-equilibrium with $\tau_{\min}=0.4$ fm. The best fit (whose parameters are shown in Fig. 1) outperforms these two scenarios, although differences are not very significant. The extracted bare jet transport coefficient is approximately $\hat{q}_{0}/T^{3}\approx 1.7$, while after logarithmic corrections the effective value is $\hat{q}/T^{3}=\hat{q}_{0}/T^{3}\log(Q_{s}^{2}/\mu_{*}^{2})\approx 7$, where $Q_{s}^{2}=\hat{q}_{0}L$, and $\mu_{*}^{2}=m_{D}^{2}e^{-2+2\gamma_{E}}/4$. We note that this extracted $\hat{q}$ is consistent with other recent extractions using Bayesian inference in the JETSCAPE framework [106].
@@
Figures 6-7 in App. A provide predictions for jets at RHIC energies [111] and charged hadrons at LHC energies [112, 113, 85, 114] using the extracted parameters. Even though our framework was designed for jets, it also provides a good description of the charged hadron $R_{AA}$ without including them in the inference. We also see, however, that our framework underestimates hadron elliptic flow at low $p_{T}$. This failure is consistent with previous findings using pQCD energy loss [68, 69]. It has been pointed out that a rather late quenching time of around $\approx 1$ fm would resolve the disagreement [69]. Alternatively, the absence of this issue for jets signals the need for model improvements which could be more relevant for (mid-$p_{T}$) hadrons than for jets. Again, for completeness, we show predictions with the previous two limiting scenarios (early pre-eq. and no pre-eq. energy loss), resulting as before in subtle differences.
@@
[…]
Due to the relatively small path lengths involved in OO, the typical value of the coherence angle is relatively large, $\theta_{c}\approx 0.5$, and so $R=0.4$ jets are unresolved and lose energy as single color charges. Therefore, hadron and jet suppression in our framework can be understood by the mere $p_{T}$ shift resulting from hadronization. In contrast, in larger systems (with smaller values of $\theta_{c}$) both $R_{AA}$ and $v_{2}$ receive additional contributions due to resolved jet substructure fluctuations [129, 44].
@@ IV Conclusions
In this work, we have integrated our semi-analytical jet quenching model with a state-of-the-art description of heavy-ion collisions, IP-Glasma+JIMWLK+MUSIC+UrQMD. We incorporated energy loss in the pre-equilibrium phase via the hydrodynamic attractor for the first time. Our framework provides a simultaneous description of jet $R_{AA}$ and $v_{2}$ in agreement with data in $AA$ collisions, offering a strong validation of the approach. We show that a joint description of these observables places constraints on the onset time of jet-medium interactions, which is around 0.2 fm, deep in the pre-equilibrium phase. Then, we applied these parameters to predict jet and hadron quenching in other collision systems. We saw that our framework successfully predicts hadron suppression but underestimates hadron elliptic flow towards lower $p_{T}$. Our results provide a complementary perspective to previous phenomenological studies. By incorporating early-time energy loss, we achieve a simultaneous description of jet $R_{AA}$ and $v_{2}$, as well as hadron $R_{AA}$ and $v_{2}$ at high-$p_{T}$. This resolves the, in our opinion, unphysical delay of quenching invoked in earlier approaches (e.g. [68, 69]) to reproduce hadron $R_{AA}$ and $v_{2}$. The successful description of jet observables supports the onset of quenching at early times, whereas the remaining discrepancy in hadron $v_{2}$ at moderate transverse momentum indicates the necessity of additional medium effects beyond longitudinal energy loss.
@@
Our constrained framework has been used to make realistic predictions for the intermediate-sized OO collisions. We obtained sizable energy loss both for jets and charged hadrons, in quantitative agreement with the results from CMS [130], where our predictions are also presented. Even though we have shown that pre-equilibrium effects play a moderate role in our model, we anticipate that including forthcoming data on high-$p_{T}$ flow coefficients will result in further constraints. Furthermore, in this work we have provided predictions for elliptic anisotropies and different centrality selections, both for hadrons and jets. Due to the small path lengths in OO, unlike in PbPb or AuAu, jet constituents are barely resolved (jets are fully coherent), and thus we obtain similar energy loss for hadrons and jets (modulo a $p_{T}$ shift in fragmentation). This outcome constitutes a prediction of our framework which can be confronted with future experimental results on jet suppression in light ions. All of our data and plots are available at [131].
@@
As an outlook, our jet quenching framework could benefit from an improved description of medium-induced emissions, elastic scatterings, and medium response in several ways. It is important to note that, at very early times, local color fields are saturated and highly anisotropic, a scenario which our energy-loss framework, built on local equilibrium, cannot accommodate. For these initial stages, it would be desirable to incorporate recent developments in energy-loss calculations for anisotropic and flowing media [132, 133, 134, 135, 136, 137, 138], correlated color fields [139], and out of-equilibrium plasma [140, 141], so as to extend quenching for those earliest times where the attractor solution is not meant to apply.